# Efficient Automated Circuit Discovery in Transformers using Contextual Decomposition

### v2

@@ 6 Discussion and Limitations
Despite having shown both qualitative and quantitative evidence of the strengths of CD-T in automated circuit discovery, our results is limited in scale and to the datasets and prior interpretation methods evaluated. More work is needed to generalize these findings to a broader set of models, datasets, and interpretation methods. Although our methods can be applied to construct circuits comprised of any of a neural network’s internal components, we only discussed and interpreted circuits built with purely attention heads. Automating the discovery of circuits with different/heterogeneous components, particularly the MLPs of a transformer, is a promising direction for further investigation. Finally, our circuit discovery is itself limited in three major ways: first, as mentioned in 4.3, the metric which measures the relevance of a source node to a target node still lacks certain properties (e.g, its scale does not correspond to an interpretable quantity, and it does not measure the sign of its contribution) that would be broadly useful in an interpretability context. Second, although the algorithm is capable of finding circuits with reasonable default settings, some amount of effort is still applied to tune the heuristic elements, such as the number of heads to select in each iteration, and the method of pruning away negative heads. Third, and more fundamentally, although our method finds circuits, it offers relatively little in the way of distilling an interpretation of the circuits which is compatible with human understanding.

### v3

@@ 5 Discussion and Limitations
Despite having shown both qualitative and quantitative evidence of the strengths of CD-T in automated circuit discovery, our results is limited in scale and to the datasets and prior interpretation methods evaluated. More work is needed to generalize these findings to a broader set of models, datasets, and interpretation methods. Although our methods can be applied to construct circuits comprised of any of a neural network’s internal components, we only discussed circuits built with purely attention heads. Automating the discovery of circuits with different/heterogeneous components, particularly the MLPs of a transformer, is a promising direction for further investigation. Finally, our method is itself limited in two ways: first, as mentioned in 3.3, the metric which measures the relevance of a source node to a target node does not differentiate signs of the contribution, which might be useful in an interpretability context. Second, while the algorithm is capable of finding circuits under reasonable default settings, some amount of effort is still applied to tune the heuristic elements, such as the number of heads to select in each iteration.
@@ A.1.3 Docstring (Heimersheim & Janiak, 2023)
Appendix B Algorithm details
B.1 Heuristics
In addition to the greedy pruning presented in Algorithm 1 as a refinement step, here we describe other heuristic elements to the algorithm.
  • The threshold for determining which set S of highest-contributing nodes may be varied: it is possible to pick the top $N$ nodes or fraction of nodes, or to automatically detect outliers. Empirically, we find that the distribution of node contributions varies quite severely, so finding a heuristic which works in all cases is actually an object of future work. In this paper, we set the threshold by varying the percentile of top nodes to extract, in the range of [90, 99] to obtain the ROC AUC in section 4.2.1.
  • Empirically we find, for a fixed target node, the relevance scores of source nodes in different layers to this target node may have different expected magnitudes, due to the numerical effects of propagation through the network. To account for this, we normalize the relevance scores by dividing by the average magnitude across scores found in a given layer.
  • To avoid the risk of propagation equations becoming numerically unstable after a large number of iterations, especially if the relevant and irrelevant constituents differ in sign at a specific index, we set the value of one of rel/irrel at this position to 0 and the other to the sum of the two terms.
B.2 Complexity analysis
To provide more clarity, the computational complexity of the algorithm satisfy the following properties:
  • Each decomposition (of a set of target nodes with respect to a set of source nodes) requires cost in FLOPs similar to one forward pass of the model (and often less, since values prior to the source nodes can be cached and values after the target nodes do not need to be calculated).
  • The core of the algorithm is a loop, where in each iteration, we search over all nodes which can potentially have high relevance to the target nodes. (This means that heuristically excluding some nodes from the search can potentially significantly decrease cost in FLOPs, and cost in FLOPs increases linearly with respect to the granularity with which we separate the nodes in the model.)
  • The memory footprint of a single decomposition, including the forward pass, is a small (less than 3) constant multiple of the cost of a forward pass; the only added costs are to keep track of the relevant and irrelevant constituent tensors separately, as well as bookkeeping of components of the target decomposition metric.
  • The cost (in FLOPs or memory) of performing the analysis on a set of input examples is linear in the number of input examples, since the same set of computations needs to be done with respect to each example.